# Some remarks on real algebraic maps which are topologically special generic maps and generalize the canonical projections of the unit spheres

### v5

@@ Abstract.
Morse functions with exactly two singular points on spheres and canonical projections of spheres are generalized to special generic maps. A special generic map is, roughly, a smooth map represented as the composition of a surjection onto a manifold with an immersion of codimension 00. Preimages of this surjections are diffeomorphic to a unit sphere in the interior of the manifold and single point sets on the boundary.
We construct real algebraic maps topologically seen as special generic maps whose images are smoothly embedded manifolds. Canonical projections of unit spheres are generalized here. We are interested in explicit and meaningful smooth maps in differential topology and real algebraic geometry. Obtaining them explicitly is important and difficult. In the stories, we previously constructed some real algebraic maps topologically special generic maps. We construct such maps in more general situations and give additional remarks.
@@
In theory of Morse functions, related singularity theory and applications to geometry of manifolds, Morse functions with exactly two singular points on spheres (in Reeb’s theorem) are of most fundamental functions and maps. Such maps are generalized to special generic maps: they are smooth maps locally represented as projections or of the form $(x_{1},\cdots x_{m})\mapsto(x_{1},\cdots x_{n-1},{\Sigma}_{j=1}^{m-n+1}{x_{j}}^{2})$. [5, 9, 44] are related pioneering studies. For Morse functions, see [41] for a pioneering study and [36] for systematic exposition on fundamental notions and arguments. For higher dimensional versions of theory of Morse functions, [52, 55] are pionnering studies.
This paper also studies real algebraic geometry. Complex algebraic geometry is presenting one of central topics in algebraic geometry. Real algebraic geometry has another nice history. [39, 53] show classical and important studies and [2, 3, 21, 22] present fundamental notions and arguments and advanced studies, for example. Our paper studies explicit construction in real algebraic geometry, respecting and generalizing (the canonical projections of) spheres.
@@
For example, [25, 40, 51] show explicit cases and some fundamental properties of real algebraic functions in the theory of Lie groups and symmetric spaces. (Unit) spheres are simplest symmetric spaces. We consider another generalization for projections of spheres: constructing real algebraic maps topologically special generic maps. The author has succeeded in such construction in [15], followed by [18] by the author himself. We present new related result. More precisely, we construct new explicit maps on some real algebraic hypersurfaces. As one of related future work, [10, 54] present closely related long-standing studies on construction of real algebraic curves in the plane and general real algebraic hypersurfaces (Remark 5).
@@
Let $X$ be a topological space with the structure of a cell complex whose dimension is finite. We can define the dimension $\dim X$ uniquely, which is a non-negative integer. A topological manifold is well-known to be homeomorphic to some CW complex. A smooth manifold is well-known to be a polyhedron canonically, which is a so-called PL manifold. It is also well-known that a topological space homeomorphic to a polyhedron whose dimension is at most $2$ has the structure of a polyhedron uniquely. If we restrict this to the class of topological manifolds, then the condition is weaken to be that the dimension is at most $3$. Consult [38] for example. Hereafter, for a manifold $X$, let $\partial X$ denote the boundary of it and ${\rm Int}\ X:=X-\partial X$ the interior of it.
@@
Let ${\mathbb{R}}^{k}$ denote the $k$-dimensional Euclidean space, which is one of most fundamental $k$-dimensional smooth manifolds. It is also the Riemannian manifold with the standard Euclidean metric. Let $\mathbb{R}:={\mathbb{R}}^{1}$. For $x\in{\mathbb{R}}^{k}$, $||x||\geq 0$ is defined as the canonically defined distance between $x$ and the origin 00. This is also regarded as a (non-singular) real algebraic manifold or the $k$-dimensional real affine space. $S^{k}:=\{x\in{\mathbb{R}}^{k+1}\mid||x||=1\}$ denotes the $k$-dimensional unit sphere, which is a $k$-dimensional smooth compact submanifold of ${\mathbb{R}}^{k+1}$ and has no boundary. It is connected for any positive integer $k\geq 1$. It is a discrete set with exactly two points for $k=0$. It is the zero set of the real polynomial $||x||^{2}-1={\Sigma}_{j=1}^{k+1}{x_{j}}^{2}-1$ ($x:=(x_{1},\cdots,x_{k+1})$) and a (non-singular) real algebraic (sub)manifold (hypersurface). $D^{k}:=\{x\in{\mathbb{R}}^{k}\mid||x||\leq 1\}$ is the $k$-dimensional unit disk. It is a $k$-dimensional smooth compact and connected submanifold of ${\mathbb{R}}^{k}$.
Let $c:X\rightarrow Y$ be a differentiable map. $x\in X$ is a singular point of $c$ if the rank of the differential at $x$ is smaller than the minimum of $\{\dim X,\dim Y\}$. $c(x)$ is called a singular value of $c$ for a singular point $x$ of $c$. Let $S(c)$ denote the set of all singular points of $c$. Here, most of differentiable maps are of the class $C^{\infty}$, in other words, smooth. A canonical projection of the Euclidean space ${\mathbb{R}}^{k}$ is denoted by ${\pi}_{k,k_{1}}:{\mathbb{R}}^{k}\rightarrow{\mathbb{R}}^{k_{1}}$. This is the map mapping each point $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}}\times{\mathbb{R}}^{k_{2}}={\mathbb{R}}^{k}$ to the first component $x_{1}\in{\mathbb{R}}^{k_{1}}$ with the conditions $k_{1},k_{2}>0$ and $k=k_{1}+k_{2}$. The canonical projection of the unit sphere $S^{k-1}$ is its restriction.
@@
Here, a real algebraic manifold means a union of connected components of the zero sets of a real polynomial map. Non-singular real algebraic manifolds are defined naturally by the implicit function theorem for the polynomial maps. The real affine space and the unit sphere give simplest examples. Our real algebraic maps mean the compositions of the canonical embeddings into the real affine spaces with the canonical projections. We may say Nash manifolds and maps for example for such maps. [46] shows related fundamental theory, arguments and some advanced theory.
@@
Spaces of smooth maps between two manifolds are topologized with so-called $C^{\infty}$ Whitney topologies. In short two smooth maps are close if their differentials of arbitrary degrees are close. $C^{r}$ Whitney topologies are also considered as weaker topologies for each non-negative integer $r$. They respect differentials of degrees at most $r$. The $C^{r_{1}}$ Whitney topologies are weaker than the $C^{r_{2}}$ Whitney topologies for $0\leq r_{1}<r_{2}$: $r_{1}$ is a non-negative integer and $r_{2}$ is a positive integer or $\infty$. Such topologies are important in singularity theory and differential topology. We omit rigorous exposition. See [11] for systematic exposition. Diffeomorphisms are smooth maps with no singular points (of them) and being homeomorphisms. The diffeomorphism group of a smooth manifold means the group of all diffeomorphisms from the manifold onto itself. It is also topologized with the $C^{\infty}$ Whitney topology as before. A smooth bundle means a bundle whose fiber is a smooth manifold and whose structure group is (a subgroup of) the diffeomorphism group. A linear bundle means a smooth bundle whose fiber is a Euclidean space, a unit disk, or a unit sphere and whose structure group consists of linear transformations, defined naturally.
We present a fundamental theorem on structures of special generic maps as Proposition 1. Using the notation, we present our new real algebraic maps as Main Theorem 1. This is proven in the next section with our additional result Main Theorems 2 and 3. Main Theorem 2 is a kind of remarks on Main Theorem 1. Main Theorem 3 is a more generalized result. The third section is on meanings of our achievement and related topics.
@@
Let $m>n\geq 1$ be integers. For a special generic map $f:M\rightarrow N$ on an $m$-dimensional closed manifold $M$ into an $n$-dimensional smooth manifold $N$ with no boundary, we have a decomposition $f=\bar{f}\circ q_{f}$ as follows.
@@
    There exist an $n$-dimensional smooth compact manifold $W_{f}$, a smooth surjection $q_{f}:M\rightarrow W_{f}$ and a smooth immersion $\bar{f}$ and $q_{f}$ maps $S(f)$ onto $\partial W_{f}$ as a diffeomorphism.
@@
    There exist a small collar neighborhood $N(\partial W_{f})$ of $\partial W_{f}$ and a linear bundle whose projection is defined as the composition of the restriction of $q_{f}$ to the preimage of $N(\partial W_{f})$ with the canonical projection to $\partial W_{f}$ and whose fiber is $D^{m-n+1}$. By (the restriction of) $q_{f}$, we also have a smooth bundle over $W_{f}-{\rm Int}\ N(\partial W_{f})$ whose fiber is the unit sphere.
@@
Let $m\geq n\geq 1$ be integers. Suppose that a smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n}}:\bar{N}\rightarrow{\mathbb{R}}^{n}$ from an $n$-dimensional compact and connected manifold $\bar{N}$ into ${\mathbb{R}}^{n}$ exists. Then there exist a pair $(M,M_{0})$ of two mutually diffeomorphic $m$-dimensional closed and connected manifolds and a pair $(f:M\rightarrow{\mathbb{R}}^{n},f_{0}:M_{0}\rightarrow{\mathbb{R}}^{n})$ of maps there and the following properties are enjoyed.
@@
    The map $f$ is special generic and in Proposition 1, $\bar{f}={\bar{f}}_{{\mathbb{R}}^{n}}:W_{f}=\bar{N}\rightarrow{\mathbb{R}}^{n}$. The map $f_{0}$ is a smooth real algebraic map on a non-singular real algebraic manifold $M_{0}\subset{\mathbb{R}}^{m+1}$ being also the zero set of a real polynomial.
@@
    There exists a pair $(\Phi:M\rightarrow M_{0},\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n})$ of homeomorphisms enjoying $\phi\circ f_{0}\circ\Phi=f$ with the two relations $\Phi(S(f))=S(f_{0})$ and $\phi(f_{0}(S_{0}(f_{0})))=f(S(f))$.
This paper is regarded as an improved version of [19] and based on our new discovery and idea.
@@
This work has been supported by the Research Institute of Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University: the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)”. The author works at Institute of Mathematics for Industry (https://www.jgmi.kyushu-u.ac.jp/en/about/young-mentors/). This is closely related to our study. Our study thanks this project for supports. The author is also a researcher at Osaka Central Advanced Mathematical Institute (OCAMI researcher), supported by MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165. He is not employed there. However this helps our studies. Our study also thanks this.
@@
Data essentially supporting our present study are all here.
@@
We also review a theorem on reconstructing special generic maps from given smooth immersions of codimension 00. This is a main theorem of [44]. In [20], followed by [17], the author has redefined several notions and consider specific reconstruction. We also abuse these notions for example.
@@
Let $m\geq n\geq 1$ be integers. Suppose that a smooth immersion ${\bar{f}}_{N}:\bar{N}\rightarrow N$ from an $n$-dimensional compact manifold $\bar{N}$ into an $n$-dimensional smooth manifold $N$ with no boundary and the following additional data (objects) are given. Hereafter we can abuse the notation in Proposition 1 and do. We can also abuse this in the case $m=n$ and we do.
  • Consider a small collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ and a smooth bundle ${\bar{M}}_{\rm Interior}={\bar{M}}_{\rm I}$ over $\bar{N}-{\rm Int}\ N(\partial\bar{N})$ whose fiber is the unit sphere $S^{m-n}\subset\partial D^{m-n+1}$. We call this bundle an interior unit sphere bundle for ${\bar{f}}_{N}$ and this is the first object.
  • A linear bundle ${\bar{M}}_{\rm Boundary}={\bar{M}}_{\rm B}$ over $\partial\bar{N}$ whose fiber is the unit disk $D^{m-n+1}$. We call this bundle a boundary linear bundle for ${\bar{f}}_{N}$ and this is the second object.
  • We consider the smooth bundle $\partial{\bar{M}}_{\rm I}$ as the restriction to the boundary $\partial(\bar{N}-{\rm Int}\ N(\partial\bar{N}))$. We also consider the subbundle $\partial{\bar{M}}_{\rm B}$ whose fiber is $S^{m-n}=\partial D^{m-n+1}$. We consider the identification between their base spaces in the natural way. More precisely, we can regard the collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ as the product $(\partial\bar{N})\times[0,1]$. Furthermore, each point $x\in\partial\bar{N}$ is identified with a point $(x,0)\in(\partial\bar{N})\times\{0\}$. In addition, the manifold $\partial N(\partial\bar{N})-\partial\bar{N}$ is identified with the space $(\partial\bar{N})\times\{1\}$ naturally and it is the subspace of $\bar{N}-{\rm Int}\ N(\partial\bar{N})\subset\bar{N}$. The third object is a bundle isomorphism $i_{\partial{\bar{M}}_{\rm I},\partial{\bar{M}}_{\rm B}}$ between the smooth bundle $\partial{\bar{M}}_{\rm I}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$. We call it an identification bundle isomorphism for ${\bar{f}}_{N}$.
Then there exist an $m$-dimensional closed manifold $M$ and a special generic map $f:M\rightarrow N$ such that $\bar{f}$ in Proposition 1 and ${\bar{f}}_{N}$ agree. Furthermore, we have the manifold and the map enjoying the following properties.
@@
    The interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ and the bundle over $\bar{N}-{\rm Int}\ N(\partial\bar{N})=W_{f}-{\rm Int}\ N(\partial W_{f})$ in Proposition 1 are isomorphic and naturally identified.
@@
    The linear bundle whose fiber is $D^{m-n+1}$ in Proposition 1 and the boundary linear bundle ${\bar{M}}_{\rm B}$ for ${\bar{f}}_{N}$ are isomorphic and naturally identified.
@@
    We also consider the restriction of the smooth bundle in (1) here to the boundary and the subbundle of the linear bundle whose fiber is $\partial D^{m-n+1}$ in (2) here. We consider their natural identification. This coincides with the identification bundle isomorphism $i_{\partial{\bar{M}}_{\rm I},\partial{\bar{M}}_{\rm B}}$ for ${\bar{f}}_{N}$.
Note that in Proposition 1, we cannot set ”$m\geq n$”. The case with the constraint $m=n$ is closely related to Eliashberg’s celebrated theory [6]. See also [7].

### v6

# Some remark on real algebraic maps which are topologically special generic maps and generalize the canonical projections of the unit spheres

@@ Abstract.
Morse functions with exactly two singular points on spheres and projections of naturally embedded spheres in the Euclidean spaces (the canonical projections of the unit spheres) are generalized to special generic maps. This is a smooth map which is the composition of a surjection onto a manifold with a smooth immersion of codimension 00. The preimage of the surjection is diffeomorphic to the unit sphere of a suitable dimension at each point of the interior of the manifold and a single point set at each point on the boundary.
We construct real algebraic maps topologically seen as special generic maps whose images are smoothly embedded manifolds. The canonical projections of the unit spheres are generalized here. We also constructed some real algebraic maps topologically special generic maps previously. We construct such maps in more general situations and give additional remarks. We are also interested in explicit and meaningful smooth maps in differential topology and real algebraic geometry. The construction is important and difficult.
@@
In theory of Morse functions, related singularity theory and applications to geometry of manifolds, Morse functions with exactly two singular points on spheres (in Reeb’s theorem) are of most fundamental functions and maps. Such maps are generalized to special generic maps: a special generic map $c$ is a smooth map locally represented as a projection or by the form $c(x_{1},\cdots x_{m})=(x_{1},\cdots x_{n-1},{\Sigma}_{j=1}^{m-n+1}{x_{j}}^{2})$. [5, 9, 44] are related pioneering studies. For Morse functions, see [41] for a pioneering study and [36] for systematic exposition on fundamental notions and arguments.
This paper also studies real algebraic geometry. Complex algebraic geometry is presenting one of central topics in algebraic geometry. Real algebraic geometry has another nice history. [39, 53] show classical and important studies and [2, 3, 21, 22] present fundamental notions and arguments and advanced studies, for example. Our paper studies explicit construction in real algebraic geometry, respecting and generalizing (the canonical projections of the unit) spheres.
@@
For example, [25, 40, 51] show explicit cases and some fundamental properties of real algebraic functions in the theory of Lie groups and symmetric spaces. The unit spheres are simplest symmetric spaces. We consider another generalization for the projections of spheres: constructing real algebraic maps topologically special generic maps. The author has succeeded in such construction in [15], followed by [18] by the author. We present new related result. More precisely, we construct new explicit maps on some real algebraic hypersurfaces. As one of related future work, [10, 54] present closely related long-standing studies on construction of real algebraic curves in the plane and general real algebraic hypersurfaces (Remark 6).
@@
Let $X$ be a topological space with the structure of a cell complex whose dimension is finite. We can define the dimension $\dim X$ uniquely, which is a non-negative integer. A topological manifold is well-known to be homeomorphic to some CW complex. A smooth manifold is well-known to be a polyhedron canonically, and a so-called PL manifold. It is also well-known that a topological space homeomorphic to a polyhedron whose dimension is at most $2$ has the structure of a polyhedron uniquely. If we restrict this to the class of topological manifolds, then the condition is weaken to be that the dimension is at most $3$. Consult [38] for example. Hereafter, for a manifold $X$, let $\partial X$ denote the boundary of it and ${\rm Int}\ X:=X-\partial X$ the interior of it.
@@
Let ${\mathbb{R}}^{k}$ denote the $k$-dimensional Euclidean space, which is one of most fundamental $k$-dimensional smooth manifolds. It is also the Riemannian manifold with the standard Euclidean metric. Let $\mathbb{R}:={\mathbb{R}}^{1}$. For $x\in{\mathbb{R}}^{k}$, $||x||\geq 0$ is defined as the canonically defined distance between $x$ and the origin 00. This is also regarded as a (non-singular) real algebraic manifold or the $k$-dimensional real affine space. $S^{k}:=\{x\in{\mathbb{R}}^{k+1}\mid||x||=1\}$ denotes the $k$-dimensional unit sphere, which is a $k$-dimensional smooth compact submanifold of ${\mathbb{R}}^{k+1}$ and has no boundary. It is connected for any positive integer $k\geq 1$. It is a discrete set with exactly two points for $k=0$. It is the zero set of the real polynomial $||x||^{2}-1={\Sigma}_{j=1}^{k+1}{x_{j}}^{2}-1$ ($x:=(x_{1},\cdots,x_{k+1})$) and a (non-singular) real algebraic (sub)manifold (hypersurface). $D^{k}:=\{x\in{\mathbb{R}}^{k}\mid||x||\leq 1\}$ is the $k$-dimensional unit disk. It is a $k$-dimensional smooth compact and connected submanifold of ${\mathbb{R}}^{k}$. Let $c:X\rightarrow Y$ be a differentiable map between differentiable manifolds $X$ and $Y$. $x\in X$ is a singular point of $c$ if the rank of the differential at $x$ is smaller than the minimum of $\{\dim X,\dim Y\}$ and $c(x)$ is called a singular value of $c$ for a singular point $x$ of $c$. Let $S(c)$ denote the set of all singular points of $c$. Here, most of differentiable maps are smooth, equivalently, of the class $C^{\infty}$. The canonical projection of the Euclidean space ${\mathbb{R}}^{k}$ is denoted by ${\pi}_{k,k_{1}}:{\mathbb{R}}^{k}\rightarrow{\mathbb{R}}^{k_{1}}$. This maps each point $x=(x_{1},x_{2})\in{\mathbb{R}}^{k_{1}}\times{\mathbb{R}}^{k_{2}}={\mathbb{R}}^{k}$ to $x_{1}\in{\mathbb{R}}^{k_{1}}$ with $k_{1},k_{2}>0$ and $k=k_{1}+k_{2}$. The canonical projection of the unit sphere $S^{k-1}$ is its restriction, rigorously.
@@
Here, a real algebraic manifold is the union of some connected components of the zero sets of a real polynomial map. Non-singular real algebraic manifolds are defined naturally by the implicit function theorem for the polynomial maps. The real affine space and the unit sphere are of simplest examples. Our real algebraic maps are the compositions of the canonical embeddings into the real affine spaces with the canonical projections. Note that they may be so-called Nash manifolds and maps, as in [39]. However, we do not use such terminologies.
@@
Spaces of smooth maps between two manifolds are topologized with so-called $C^{\infty}$ Whitney topologies. Roughly, two smooth maps are close if their differentials of arbitrary degrees are close. $C^{r}$ Whitney topologies are weaker topologies, respecting differentials of degrees at most $r$. The $C^{r_{1}}$ Whitney topologies are weaker than the $C^{r_{2}}$ Whitney topologies for $0\leq r_{1}<r_{2}$: $r_{1}$ is a non-negative integer and $r_{2}$ is a positive integer or $\infty$. These topologies are important in singularity theory and differential topology. Rigorous exposition on singularity theory of differential maps including these topologies is left to [11]. Diffeomorphisms are smooth homeomorphisms with no singular points (of them). The diffeomorphism group of a smooth manifold is the group of all diffeomorphisms from the manifold onto itself. A smooth bundle is a bundle whose fiber is a smooth manifold and whose structure group is (a subgroup of) the diffeomorphism group. A linear bundle is a smooth bundle whose fiber is the Euclidean space, the unit disk, or the unit sphere and whose structure group consists of linear transformations, defined naturally.
Proposition 1 explains a fundamental theorem on structures of special generic maps. Using the notation, we present our new real algebraic map which is special generic topologically in Main Theorem 1, proven in the next section with our additional result Main Theorems 2 and 3. Main Theorem 2 is for additional comments on Main Theorem 1. Main Theorem 3 generalizes them. The third section presents additional comments on our new result.
@@
Let $m>n\geq 1$ be integers. For a special generic map $f:M\rightarrow N$ from an $m$-dimensional closed manifold $M$ into an $n$-dimensional smooth manifold $N$ with no boundary, we have a decomposition $f=\bar{f}\circ q_{f}$ as follows. Note that the case $N:={\mathbb{R}}^{n}$ is mainly discussed in [44] and our paper, and that as is also presented shortly in [44], $N$ can be general.
@@
    There exist an $n$-dimensional smooth compact manifold $W_{f}$, a smooth surjection $q_{f}:M\rightarrow W_{f}$, mapping $S(f)$ onto $\partial W_{f}$ by a diffeomorphism, and a smooth immersion $\bar{f}$.
@@
    There exist a small collar neighborhood $N(\partial W_{f})$ of $\partial W_{f}$ and a linear bundle whose projection is the composition of the restriction of $q_{f}$ to the preimage of $N(\partial W_{f})$ with the canonical projection to $\partial W_{f}$ and whose fiber is $D^{m-n+1}$. By (the restriction of) $q_{f}$, we also have a smooth bundle over the closure of $W_{f}-N(\partial W_{f})$ in $W_{f}$ whose fiber is the unit sphere $S^{m-n}$.
@@
Let $m\geq n\geq 1$ be integers. Suppose that a smooth embedding ${\bar{f}}_{{\mathbb{R}}^{n}}:\bar{N}\rightarrow{\mathbb{R}}^{n}$ from an $n$-dimensional compact and connected manifold $\bar{N}$ into ${\mathbb{R}}^{n}$ exists. Then we have an $m$-dimensional non-singular real algebraic manifold $M$ being also compact, connected, and the zero set of a real polynomial function and the following maps.
@@
    A special generic map $f:M\rightarrow N:={\mathbb{R}}^{n}$ in Proposition 1 with $\bar{f}={\bar{f}}_{{\mathbb{R}}^{n}}:W_{f}=\bar{N}\rightarrow{\mathbb{R}}^{n}$. This map $f$ is not real algebraic in general.
@@
    A real algebraic map $f_{0}:M\rightarrow{\mathbb{R}}^{n}$ and a homeomorphism $\phi:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$ satisfying the relation $f_{0}=\phi\circ f$ and $\phi(f(S(f)))=f_{0}(S(f_{0}))$: remember that $S(c)$ denotes the singular set of a smooth map $c$.
The simplest case $\bar{N}=D^{n}$ accounts for the canonical projection $f=f_{0}:S^{m}\rightarrow{\mathbb{R}}^{n}$ of the unit sphere $S^{m}$ with $\phi$ being the identity map.
@@
This work was also supported by the Research Institute of Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University: the conference ”Extension of the singularity theory” (https://www2.akita-nct.ac.jp/kasedou/workshop/rims2023/index.html)”. The author works at Institute of Mathematics for Industry (https://www.jgmi.kyushu-u.ac.jp/en/about/young-mentors/). This is closely related to our study. Our study thanks this project for supports. The author is also a researcher at Osaka Central Advanced Mathematical Institute (OCAMI researcher), supported by MEXT Promotion of Distinctive Joint Research Center Program JPMXP0723833165. He is not employed there. However this helps our studies. Our study also thanks this.
@@
Data essentially supporting our present study are all here. This paper is regarded as an improved version of one of a preprints of the author [19].
@@
As an important ingredient in special generic maps, we explain the Morse function for the natural height of the disk $D^{k}\subset{\mathbb{R}}^{k}$. We map $x\in D^{k}$ to $(\sqrt{1-{||x||}^{2}},x)\in\mathbb{R}\times D^{k}\subset{\mathbb{R}}^{k+1}$ by a diffeomorphism, compose the projection to the second component and we have the Morse function for the natural height of the disk $D^{k}\subset{\mathbb{R}}^{k}$. The projection of $D^{k}$ embedded in ${\mathbb{R}}^{k+1}$ here is the restriction of ${\pi}_{k+1,1}{\mid}_{S^{k}}$ to $S^{k}\bigcap\{x:=(x_{1},\cdots x_{k+1})\mid x_{1}\geq 0\}$. Locally, a special generic map from an $m$-dimensional manifold with no boundary into an $n$-dimensional manifold with no boundary with $m\geq n$ is a smooth map such that around each singular point of it, the map is represented as the product map of such a function and the identity map on an ($n-1$)-dimensional manifold (where suitable local coordinates are chosen).
Related to Proposition 1, we also review a theorem reconstructing special generic maps from given smooth immersions of codimension 00, conversely. Note again that the case $N:={\mathbb{R}}^{n}$ is discussed in the original study mainly, and that we can consider the general case. In [20], followed by [17], the author has redefined several notions and consider specific reconstruction. We also abuse these notions for example.
@@
Let $m\geq n\geq 1$ be integers. Suppose that a smooth immersion ${\bar{f}}_{N}:\bar{N}\rightarrow N$ from an $n$-dimensional compact manifold $\bar{N}$ into an $n$-dimensional smooth manifold $N$ with no boundary and the following additional data (objects) are given. Hereafter, we can abuse the notation in Proposition 1 and do. We can also abuse this in the case $m=n$ and we do.
  • A small collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ and a smooth bundle ${\bar{M}}_{\rm Interior}={\bar{M}}_{\rm I}$ over the closure $\overline{\bar{N}-N(\partial\bar{N})}$ of $\bar{N}-N(\partial\bar{N})$ in $\bar{N}$ whose fiber is the unit sphere $S^{m-n}\subset\partial D^{m-n+1}$. We call this bundle an interior unit sphere bundle for ${\bar{f}}_{N}$.
  • A linear bundle ${\bar{M}}_{\rm Boundary}={\bar{M}}_{\rm B}$ over $\partial\bar{N}$ whose fiber is the unit disk $D^{m-n+1}$. We call this bundle a boundary linear bundle for ${\bar{f}}_{N}$.
  • We consider the smooth bundle $\partial{\bar{M}}_{\rm I}$ as the restriction to the boundary $\partial(\overline{\bar{N}-N(\partial\bar{N})})$ of the closure $\overline{\bar{N}-N(\partial\bar{N})}$. We also consider the subbundle $\partial{\bar{M}}_{\rm B}$ whose fiber is $S^{m-n}=\partial D^{m-n+1}$. We consider the identification between their base spaces in the natural way. More precisely, we can regard the collar neighborhood $N(\partial\bar{N})$ of $\partial\bar{N}$ as the product $(\partial\bar{N})\times[0,1]$. Furthermore, each point $x\in\partial\bar{N}$ is naturally identified with a point $(x,0)\in(\partial\bar{N})\times\{0\}$, of course. In addition, in considering the base space of ${\bar{M}}_{\rm Boundary}={\bar{M}}_{\rm B}$ over $\partial N(\partial\bar{N})-\partial\bar{N}=(\partial\bar{N})\times\{1\}$ as $\partial\bar{N}$, $x\in\partial\bar{N}$ and $(x,0)\in(\partial\bar{N})\times\{0\}$ are also seen as $(x,1)\in(\partial\bar{N})\times\{1\}$.
    In addition, the manifold $\partial N(\partial\bar{N})-\partial\bar{N}=(\partial\bar{N})\times\{1\}$ is also regarded as the subspace of $\overline{\bar{N}-N(\partial\bar{N})}\subset\bar{N}$ and these two spaces are naturally identified. The third object is a bundle isomorphism between the smooth bundle $\partial{\bar{M}}_{\rm I}$ over $\partial(\overline{\bar{N}-N(\partial\bar{N})})\subset\overline{\bar{N}-N(\partial\bar{N})}$ and the smooth bundle $\partial{\bar{M}}_{\rm B}$ over $\partial\bar{N}=\partial N(\partial\bar{N})-\partial\bar{N}$ where the induced map between the base spaces is the natural identification for the spaces of course: note also that we identify several spaces according to the identification before. We call this bundle isomorphism an identification bundle isomorphism for ${\bar{f}}_{N}$.
Then there exist an $m$-dimensional closed manifold $M$ and a special generic map $f:M\rightarrow N$ such that $\bar{f}$ in Proposition 1 and ${\bar{f}}_{N}$ agree. We can also have the manifold and the map enjoying the following properties.
@@
    The interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ and the bundle over $\overline{\bar{N}-N(\partial\bar{N})}=\overline{W_{f}-N(\partial W_{f})}$ in Proposition 1 are isomorphic and naturally identified via a diffeomorphism ${\Phi}_{M,{\rm I}}:{\bar{M}}_{\rm I}\rightarrow M-{q_{f}}^{-1}(N(\partial\bar{N}))$ and the identity map on $\overline{\bar{N}-N(\partial\bar{N})}=\overline{W_{f}-N(\partial W_{f})}$ where the closure is taken in $\bar{N}=W_{f}$ as before.
@@
    The linear bundle whose fiber is $D^{m-n+1}$ in Proposition 1 and the boundary linear bundle ${\bar{M}}_{\rm B}$ for ${\bar{f}}_{N}$ are isomorphic and naturally identified via a diffeomorphism ${\Phi}_{M,{\rm B}}:{\bar{M}}_{\rm B}\rightarrow{q_{f}}^{-1}(N(\partial\bar{N}))$ and the identity map on $N(\partial\bar{N})=N(\partial W_{f})$.
@@
    We consider the bundle over $\overline{\bar{N}-N(\partial\bar{N})}=\overline{W_{f}-N(\partial W_{f})}$ and the linear bundle whose fiber is $D^{m-n+1}$ in Proposition 1. Their restrictions to the boundaries are identified by the identification induced canonically from the manifold $M$ and the map $f:M\rightarrow N$.
    We also consider the restriction of the smooth bundle ${\bar{M}}_{\rm I}$ in (1) to the boundary and the subbundle of the linear bundle ${\bar{M}}_{\rm B}$ whose fiber is $\partial D^{m-n+1}$ in (2). We consider their identification or the bundle isomorphism induced from (1, 2) and the short arguments on $M$ and $f:M\rightarrow N$ just before: the identification coincides with the identification bundle isomorphism for ${\bar{f}}_{N}$.
Remark 1.
This remarks reconstruction of a special generic map $f:M\rightarrow N$ in Proposition 2.
  1. (1)
    By considering the identification bundle isomorphism for ${\bar{f}}_{N}$ and gluing (the spaces of the domains of) ${\Phi}_{M,{\rm I}}$ and ${\Phi}_{M,{\rm B}}$ by this, we have a diffeomorphism ${\Phi}_{M,{\rm I},{\rm B}}$ onto our new manifold $M$ such that ${\Phi}_{M,{\rm I}}$ and ${\Phi}_{M,{\rm B}}$ are regarded as its restrictions.
  2. (2)
    We can regard that the interior unit sphere bundle ${\bar{M}}_{\rm I}$ for ${\bar{f}}_{N}$ and the boundary linear bundle ${\bar{M}}_{\rm B}$ for ${\bar{f}}_{N}$ naturally reconstructs a smooth surjection $q_{f}:M\rightarrow W_{f}=\bar{N}$. After that, we compose the given smooth immersion ${\bar{f}}_{N}:\bar{N}\rightarrow N$ to have our desired special generic map $f:M\rightarrow N$.
Note that in Proposition 1, we cannot consider the condition ”$m\geq n$”. The case with the constraint $m=n$ is closely related to Eliashberg’s celebrated theory [6]. See also [7].